Can someone please help explain?
The minute hand of a clock is 11 inches long and moves from 12 to 4 o'clock. How far does the tip of the minute hand move? Express your answer in terms of pi and then round two decimal places
thank you!

Answer by TimothyLamb(4379)

on the clock face 12 = 0
---
number of segments in the arc from 12 to 4?
number of segments = 4 - 12
but 12 = 0
number of segments = 4 - 0 = 4
---
the clock face is divided into 12 segments (5 minutes each)
---
what fraction of the clock circle is the arc from 12 to 4?
4/12 = 1/3
---
in one hour the tip of the minute hand traces an arc of length c = 2*pi*r
where r = 11 inches
c = 2*pi*11 = 22 * pi
---
what fraction of c is the arc from 12 to 4?
1/3 * 22 * pi
---
answer:
7.33 * pi
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php